solve the equation for y
10x-5y=25

Answer by ewatrrr(24785)


Hi,
solving the equation for y
10x-5y=25
10x - 25 = 5y
2x - 5 = y
y = 2x - 5 Slope = 2 Or (2/1)(positive.. line slants to the right)
Pt(0,-5) and Pt(5/2, 0) points on the line
Note: as y increases/decreases by 2, then x increases/decreases by 1